# Magnetic Field Amplification and Particle Acceleration in Weakly Magnetized Trans-relativistic Electron-ion Shocks

### v1

@@ Abstract
We investigate the physics of quasi-parallel trans-relativistic shocks propagating in weakly magnetized plasmas by means of long-duration two-dimensional particle-in-cell simulations. The structure of the shock precursor is shaped by a competition between the Bell instability and the Weibel instability. The Bell instability is dominant at relatively high magnetizations $(\sigma\gtrsim 10^{-3})$, whereas the Weibel instability prevails at lower magnetizations $(\sigma\lesssim 10^{-4})$. Bell-dominated shocks efficiently accelerate ions, converting a fraction $\varepsilon_{\mathrm{i}}\sim 0.2$ of the upstream flow energy into downstream nonthermal ion energy. The maximum energy of nonthermal ions exhibits a Bohm scaling in time, as $E_{\max}\propto t$. A much smaller fraction $\varepsilon_{\mathrm{e}}\ll 0.1$ of the upstream flow energy goes into downstream nonthermal electrons in the Bell-dominated regime. On the other hand, Weibel-dominated shocks efficiently generate both nonthermal ions and electrons with $\varepsilon_{\mathrm{i}}\sim\varepsilon_{\mathrm{e}}\sim 0.1$, albeit with a slower scaling for the maximum energy, $E_{\mathrm{max}}\propto t^{1/2}$. Our results are applicable to a wide range of trans-relativistic shocks, including the termination shocks of extragalactic jets, the late stages of gamma-ray burst afterglows, and shocks in fast blue optical transients.
@@ 1 Introduction
Collisionless shocks are among the most efficient particle accelerators in the universe (Drury1983; Blandford1987). They can be sources of cosmic rays and intense nonthermal emission. Ultra-relativistic shocks, with shock Lorentz factors $\Gamma_{\mathrm{sh}}\gg 10$, have been studied extensively in the context of gamma-ray burst (GRB) afterglow emission and ultra-high-energy cosmic ray (UHECR) production. Theory and particle-in-cell (PIC) simulations have revealed that the Weibel filamentation instability (Weibel1959; Fried1959; Medvedev1999; Silva2003) can generate intense magnetic fields in these shocks. PIC simulations also show efficient particle acceleration by the Fermi process (Spitkovsky2008b; Sironi2013; Groselj2024).
@@
Late-time observations of the binary neutron star merger event GW170817 (Margutti2018; Hajela2019; Hajela2022) and the recent discovery of fast blue optical transients (FBOTs), such as the AT2018cow (Ho2019; Margutti2019), direct our attention to trans-relativistic shocks, with Lorentz factors of a few. The most important difference between trans-relativistic shocks and ultra-relativistic shocks is that trans-relativistic shocks can generally be subluminal, i.e., high-energy particles can outrun the shock and escape far upstream along magnetic field lines (Sironi2009). This allows a potential amplification of the upstream magnetic field via the Bell streaming instability, similar to non-relativistic shocks (Bell2004; Caprioli2014b; Park2015). At the same time, the mean energy per particle of trans-relativistic shocks is much larger than their non-relativistic counterparts, which makes them compelling candidates for UHECR production and bright synchrotron emission.
@@
Crumley2019 performed PIC simulations of trans-relativistic shocks with $\sigma\gtrsim 10^{-3}$, where $\sigma$ is the ratio of the ambient magnetic field energy to the plasma rest-mass energy (see Section 2). They show that Bell instability can indeed be triggered in the upstream of trans-relativistic shocks, and those shocks are efficient hadronic accelerators. Their results are applicable to trans-relativistic shocks with moderate upstream magnetization, e.g., the termination shocks of active galactic nuclei (AGN) jets and microquasars. However, late-time GRB afterglows and FBOTs propagate into weakly magnetized media, e.g., a typical interstellar-like medium has a magnetization of $\sigma\sim 10^{-9}$. The shock micro-physics in this parameter range could be qualitatively different from that at higher magnetizations. Most notably, the fraction of energy transferred to nonthermal electrons, in Crumley2019, seems too small to explain the bright X-ray emission from the trans-relativistic stage of GW170817 (Hajela2019).
@@
In this paper, we investigate the physics of weakly magnetized trans-relativistic shocks with long-duration PIC simulations. We show that there is a transition from a Bell-dominated regime at high magnetizations $(\sigma\gtrsim 10^{-3})$, to a Weibel-dominated regime at low magnetizations $(\sigma\lesssim 10^{-4})$. Ions are accelerated more efficiently, as compared to electrons, in Bell-dominated cases. On the other hand, Weibel-dominated cases exhibit significantly larger nonthermal electron energy fractions. Our results are primarily applicable to the late stages of GRB afterglows and FBOTs, but we also comment on the implications of our work for the termination shocks of relativistic jets.
2 Method
@@
We perform 2D PIC simulations with the OSIRIS code (Fonseca2002; Fonseca2013). Figure 1 shows the schematics of the configuration. Panel (a) corresponds to the upstream frame, in which the following physical quantities that characterize a collisionless shock are defined. The shock velocity $V_{\mathrm{sh}}$ is the speed at which the shock front moves in the upstream. We set up the simulations so that the expected shock Lorentz factor in the upstream frame is $\Gamma_{\mathrm{sh}}=1/\sqrt{1-(V_{\mathrm{sh}}/c)^{2}}\sim 2$. The angle $\theta_{B}$ is the angle between the ambient magnetic field direction and the shock normal. We focus on quasi-parallel shocks with $\theta_{B}=20^{\circ}$. Finally, the shock magnetization $\sigma$ is defined here in the upstream frame as:
@@
We made the following optimizations to minimize any numerical artifacts. An alternative stencil to the standard Yee stencil (Blinne2018) is used to suppress the numerical Cherenkov instability (see Groselj2022). We apply 16 passes, in each direction, of binomial filtering to the electric current at each time step to further reduce the noise. These strategies allowed us to evolve the system for an unprecedentedly long time of $\omega_{\mathrm{pi}}t=7000$, and up to $\omega_{\mathrm{pi}}t=12000$ in some cases. We initialize the upstream plasma with a velocity profile that is zero at the $x=0$ boundary, and linearly transitions to the upstream value $V_{0}$ at $x=100\,d_{\mathrm{i}}$ (Groselj2024). The choice of initial flow profile suppresses the unphysically strong initial reflection of particles from the left wall before the shock is formed.
3 Results
3.1 Shock Structure
@@
Interaction between the cosmic rays and the incoming upstream plasma results in a structure typical of the late stage of the Bell instability (Caprioli2014b; Crumley2019). In this paper, we define cosmic rays as nonthermal particles that are reflected and energized at the shock and propagate into the upstream. The quantitative definition and identification method of these cosmic rays in the simulations are discussed in Subsections 3.2 and 3.3. The fluctuations, particularly in the $B_{y}$ component (panel (c)), have a wavevector predominantly parallel to the ambient field, with a typical spatial scale of a few tens of ion skin depths, as expected for the Bell instability (Bell2004; Amato2009). Locally, the fields reach near equipartition strength. The late nonlinear stage, which we are looking at, displays a large-scale structure of rarified cavities and dense filaments (panel (a)) and $B_{z}$ (panel (d)) (Caprioli2014a; Caprioli2014b). Since the Bell modes, in their linear stage, have a circular polarization, the $B_{z}$ structure is similar to $B_{y}$, modulo this nonlinear density structure.
@@
These results are consistent with Crumley2019, in which the parameters in their fiducial run correspond to $\Gamma_{\mathrm{sh}}\sim 1.8,\theta_{B}=10^{\circ}$, and $\sigma=3\times 10^{-3}$, respectively, with our definitions, which is relatively similar to the run in Figure 2. Their simulation ends at $\omega_{\mathrm{pi}}t\sim 4000$. Since we have run the simulation longer than they have, the nonlinear filamentary structure is more pronounced.
@@
Panels (e,f) are the phase space densities $x-p_{x}$ of the two species, in which $p_{x}$ is normalized by the momentum of the upstream ions $m_{\mathrm{i}}\gamma_{0}v_{0}$. We can see that there is a significant amount of high-energy ions, with $|p_{x}|/m_{\mathrm{i}}\gamma_{0}v_{0}\gg 1$, in the downstream. However, there is a much smaller number of electrons in the same energy range. This implies that the particle acceleration efficiency is different between the two species. We will discuss particle acceleration in Subsection 3.3.
@@
Figure 2: Snapshot of the $\sigma=10^{-3}$ shock, taken at $\omega_{\mathrm{pi}}t=7000$. Panel (a) is the plasma density $N_{\mathrm{e}}$ normalized by the upstream value. Panels (b-d) are the magnetic field components in units of the equipartition magnetic field, Equation (2). The colorbars are in a symmetric log scale, in which the range [0.01,0.01][-0.01,0.01] is in a linear scale, and values outside this range are in a log scale. Panels (e, f) are the phase space densities $f(x,p_{x})$ of ions and electrons, respectively.
@@
Figure 3 shows the $\sigma=10^{-3.5}$ shock at the same time in the simulation $(\omega_{\mathrm{pi}}t=7000)$. The fluctuations that dominate in the upstream $B_{y}$ component, panel (c), still have a wave vector parallel to the ambient field, similarly to the $\sigma=10^{-3}$ case, but their amplitude is much smaller. This can be attributed to the cosmic ray current satisfying the high-current condition (Weidl2019). The high-current condition is satisfied when the cosmic ray current $J_{\mathrm{CR}}$ in the upstream becomes larger than a critical value, which depends on the upstream magnetization. We will discuss this in Subsection 3.2. We can see in Figure 3 (e) that there are more reflected ions in the $\sigma=10^{-3.5}$ case, compared to the $\sigma=10^{-3}$ case (Figure 2 (e)). The difference in the number of downstream high-energy particles between ions and electrons is less significant than in the $\sigma=10^{-3}$ case.
@@
Figure 3: Snapshot of the $\sigma=10^{-3.5}$ shock, taken at $\omega_{\mathrm{pi}}t=7000$. The format is the same as in Figure 2.
@@
Finally, Figure 4 is the $\sigma=10^{-4}$ shock at $\omega_{\mathrm{pi}}t=7000$. The magnetic field structure in this case is distinct from the previous cases. Most of the fluctuating magnetic field energy is in the $B_{z}$ component, panel (d), and little in $B_{y}$. The wavenumber vector is mostly perpendicular to the ambient field, i.e., $\bm{k}\perp\hat{\bm{B}}_{0}$, with a characteristic spatial scale of $d_{\mathrm{i}}k\sim 1$ in the downstream frame. These are the characteristics of the Weibel instability (Weibel1959; Fried1959). Note that cosmic rays are still streaming along the ambient field, and the magnetic field could alter the structure of the Weibel modes as compared to the case of an unmagnetized medium (Lemoine2014; Grassi2017). We discuss the unmagnetized $\sigma=0$ case in Appendix A.
@@
Figure 4: Snapshot of the $\sigma=10^{-4}$ shock, taken at $\omega_{\mathrm{pi}}t=7000$. The format is the same as in Figure 2. Compared to Figure 2, the shock structure is notably changed, as the precursor is now shaped by the Weibel instability.
@@
Here, we have compared the structures of shocks with fixed $\Gamma_{\mathrm{sh}}\sim 2$ and $\theta_{B}=20^{\circ}$, varying the upstream $\sigma$. We have shown that the dominant physics in the upstream has a transition from the Bell instability at high magnetizations $\sigma\gtrsim 10^{-3}$ to the Weibel instability at lower magnetizations $\sigma\lesssim 10^{-4}$. In between, at $\sigma=10^{-3.5}$ we see a magnetic field structure, morphologically similar to Bell modes, but having a much smaller amplitude. In Subsection 3.2, we discuss how the magnetization and cosmic ray properties determine the dominant instability. The $\sigma=10^{-4.5}$ case is not shown here because the overall structure is essentially the same as the $\sigma=10^{-4}$ case. Nevertheless, it will be included in the quantitative analysis of the following Subsections.
3.2 Cosmic Ray Driven Instabilities
@@
Here, we discuss the plasma streaming instabilities in the upstream, driven by the returning cosmic rays. In Subsection 3.1, we discussed how the precursor is shaped by a competition between the Bell instability and the Weibel instability. It has recently been shown that the characteristics of Bell instability change when the current carried by the cosmic rays exceeds a certain threshold (Weidl2019). This is called the high-current regime. Lichko2025 suggests that the fluctuating fields $\delta B$ generated by Bell instability in the high-current regime saturate at a fraction of the ambient magnetic field energy. In their work, the typical saturation value was $\delta B/B_{0}\sim 5$. We find a similar value of $\delta B/B_{0}\sim 10$ for our parameters (see Appendix B). This is different from the classical low-current regime of Bell instability, in which a fixed fraction $\sim 0.1$ of the cosmic ray momentum flux is converted into the amplified fields (Caprioli2014b; Zacharegkas2024). The saturation energy density of the high-current Bell mode scales with $B_{0}^{2}$, while that of the low-current Bell mode scales with the cosmic ray momentum flux. It is essential to distinguish these two regimes, especially because we need to understand their competition with the Weibel instability. The high-current condition is satisfied when
@@
To investigate the nature of Bell mode saturation, we explicitly check Equation (3) in our PIC simulations. To this end, we define cosmic rays as the ions in the upstream region that have changed the sign of their $x$-momentum, at least once, i.e., that have been reflected back by the shock (Lemoine2019; Vanthieghem2022). We measure the cosmic ray ion four current $J^{\mu}_{\mathrm{CR}}$ at $x-x_{\mathrm{sh}}(t)]/d_{\mathrm{i}}=[200,300]$ in the simulation frame. Then we compute the upstream cosmic ray current along the background magnetic field as $J_{\mathrm{CR}}=J^{x}_{\mathrm{CR}}|_{\mathrm{u}}/\cos{\theta_{B}}$, where $|_{\mathrm{u}}$ indicates quantities measured in the upstream frame, and $\theta_{B}$ is defined in the upstream frame. In addition, we used the $y$ component and computed $J^{y}_{\mathrm{CR}}|_{\mathrm{u}}/\sin\theta_{B}$. The two methods give consistent results in this region. When the current is measured closer to the shock front, the two estimates are less consistent due to the change in the background plasma profile, e.g., resulting from the deceleration of the upstream flow (Vanthieghem2022). We note that we have assumed that the cosmic rays are dominated by ions. This is a common choice for Bell instability in non-relativistic shocks (Bell2004; Amato2009). We confirm that this is also valid in trans-relativistic shocks, in Subsection 3.3. The role of electron cosmic rays is discussed in Appendix C.
@@
Figure 5 shows the time evolution of the cosmic ray current averaged in the spatial interval $x-x_{\mathrm{sh}}(t)]/d_{\mathrm{i}}=[200,300]$. Panel (a) shows the raw $\eta$, whereas panel (b) corresponds to $\eta/\eta_{\mathrm{crit}}$. In panel (b), we can see that the $\sigma=10^{-3}$ shock eventually self-regulates the current to a subcritical value, consistent with the strong low-current Bell magnetic fluctuations in Figure 2. On the other hand, the cases with magnetization below $\sigma=10^{-3.5}$ remain supercritical, at least up to $\omega_{\mathrm{pi}}t=7000$. Looking at $1000\lesssim\omega_{\mathrm{pi}}t\lesssim 2000$ in panel (a), we notice that the two high-$\sigma$ cases are clearly distinct from the two low-$\sigma$ cases. This is presumably because of the difference in the ion reflection process. The classical reflection mechanism related to shock potential and mirror force (Balogh2013; Burgess2015) could apply to the high-$\sigma$ cases. On the other hand, the reflection in the low-$\sigma$ cases is governed by interaction with the small-scale Weibel fields in the vicinity of the shock front (Spitkovsky2008a; Kato2008; Jikei2024b; Parsons2024). This results in the $\sigma=10^{-3.5}$ case, Figure 3, being more supercritical, i.e., larger $\eta/\eta_{\mathrm{crit}}$, than the $\sigma=10^{-4}$ case, despite the larger $\eta_{\mathrm{crit}}$. Below $\sigma=10^{-4}$, the cosmic ray current is almost independent of $\sigma$; therefore, lower magnetization shocks would be more supercritical.

### v2

@@ Abstract
We investigate the physics of quasi-parallel trans-relativistic shocks propagating in weakly magnetized plasmas by means of long-duration two-dimensional particle-in-cell simulations. The structure of the shock precursor is shaped by a competition between the Bell instability and the Weibel (filamentation) instability. The Bell instability is dominant at relatively high magnetizations $(\sigma\gtrsim 10^{-3})$, whereas the Weibel instability prevails at lower magnetizations $(\sigma\lesssim 10^{-4})$. Shocks with precursors shaped by Bell modes efficiently accelerate ions, converting a fraction $\varepsilon_{\mathrm{i}}\sim 0.2$ of the upstream flow energy into downstream nonthermal ion energy. The maximum energy of nonthermal ions exhibits a Bohm scaling in time, as $E_{\max}\propto t$. A much smaller fraction $\varepsilon_{\mathrm{e}}\ll 0.1$ of the upstream flow energy goes into downstream nonthermal electrons in the Bell regime. On the other hand, when the precursor is dominated by Weibel modes, the shock efficiently generates both nonthermal ions and electrons with $\varepsilon_{\mathrm{i}}\sim\varepsilon_{\mathrm{e}}\sim 0.1$, albeit with a slower scaling for the maximum energy, $E_{\mathrm{max}}\propto t^{1/2}$. Our results are applicable to a wide range of trans-relativistic shocks, including the termination shocks of extragalactic jets, the late stages of gamma-ray burst afterglows, and shocks in fast blue optical transients.
@@ I Introduction
Collisionless shocks are among the most efficient particle accelerators in the universe (Drury, 1983; Blandford and Eichler, 1987). They can be sources of cosmic rays and intense nonthermal emission. Ultra-relativistic shocks, with shock Lorentz factors $\Gamma_{\mathrm{sh}}\gg 10$, have been studied extensively in the context of gamma-ray burst (GRB) afterglow emission and ultra-high-energy cosmic ray (UHECR) production. Theory and particle-in-cell (PIC) simulations have revealed that the Weibel filamentation instability (Weibel, 1959; Fried, 1959; Medvedev and Loeb, 1999; Silva et al., 2003) can generate intense magnetic fields in these shocks.  [^1] Strictly speaking, the Weibel instability (driven by temperature anisotropy) and the filamentation instability (driven by counterstreaming beams) should be treated separately (e.g., Bret, 2009). Hereafter, however, we refer to the latter as the Weibel instability, following the prevailing convention in the astrophysical community. PIC simulations also show efficient particle acceleration by the Fermi process (Spitkovsky, 2008b; Sironi et al., 2013; Grošelj et al., 2024).
@@
Late-time observations of the binary neutron star merger event GW170817 (Margutti et al., 2018; Hajela et al., 2019, 2022) and the recent discovery of fast blue optical transients (FBOTs), such as the AT2018cow (Ho et al., 2019; Margutti et al., 2019), direct our attention to trans-relativistic shocks, with Lorentz factors of a few. The most important difference between trans-relativistic shocks and ultra-relativistic shocks is that trans-relativistic shocks can generally be subluminal, i.e., high-energy particles can outrun the shock and escape far upstream along magnetic field lines (Sironi and Spitkovsky, 2009). This allows a potential amplification of the upstream magnetic field via the Bell streaming instability, similar to non-relativistic shocks (Bell, 2004; Caprioli and Spitkovsky, 2014b; Park et al., 2015). At the same time, the mean energy per particle of trans-relativistic shocks is much larger than their non-relativistic counterparts, which makes them compelling candidates for UHECR production and bright synchrotron emission.
@@
Crumley et al. (2019) performed PIC simulations of trans-relativistic shocks with $\sigma\gtrsim 10^{-3}$, where $\sigma$ is the ratio of the ambient magnetic field energy to the plasma rest-mass energy (see Section II). They show that Bell instability can indeed be triggered in the upstream of trans-relativistic shocks, and those shocks are efficient hadronic accelerators. Their results are applicable to trans-relativistic shocks with moderate upstream magnetization, e.g., the termination shocks of active galactic nuclei (AGN) jets and microquasars. However, late-time GRB afterglows and FBOTs propagate into weakly magnetized media, e.g., a typical interstellar-like medium has a magnetization of $\sigma\sim 10^{-9}$. The shock micro-physics in this parameter range could be qualitatively different from that at higher magnetizations. Most notably, the fraction of energy transferred to nonthermal electrons, in Crumley et al. (2019), seems too small to explain the bright X-ray emission from the trans-relativistic stage of GW170817 (Hajela et al., 2019).
@@
In this paper, we investigate the physics of weakly magnetized trans-relativistic shocks with long-duration PIC simulations. We show that there is a transition from a Bell-dominated regime at high magnetizations $(\sigma\gtrsim 10^{-3})$, to a Weibel-dominated regime at low magnetizations $(\sigma\lesssim 10^{-4})$. These plasma instabilities could also play a role at the shock front and in the downstream region, but we focus on those excited in the upstream region (see Section III for details). Ions are accelerated more efficiently, as compared to electrons, in Bell-dominated cases. On the other hand, Weibel-dominated cases exhibit significantly larger nonthermal electron energy fractions. Our results are primarily applicable to the late stages of GRB afterglows and FBOTs, but we also comment on the implications of our work for the termination shocks of relativistic jets.
II Method
@@
We perform 2D PIC simulations with the OSIRIS code (Fonseca et al., 2002, 2013). Figure 1 shows the schematics of the configuration. Panel (a) corresponds to the upstream frame, in which the following physical quantities that characterize a collisionless shock are defined. The shock velocity $V_{\mathrm{sh}}$ is the speed at which the shock front moves in the upstream. We set up the simulations so that the expected shock Lorentz factor in the upstream frame is $\Gamma_{\mathrm{sh}}=1/\sqrt{1-(V_{\mathrm{sh}}/c)^{2}}\sim 2$. The angle $\theta_{B}$ is the angle between the ambient magnetic field direction and the shock normal. We focus on quasi-parallel shocks with $\theta_{B}=20^{\circ}$. Finally, the shock magnetization $\sigma$ is defined here in the upstream frame as:
@@
We made the following optimizations to minimize any numerical artifacts. An alternative stencil to the standard Yee stencil (Blinne et al., 2018) is used to suppress the numerical Cherenkov instability (see Grošelj et al., 2022). We apply 16 passes, in each direction, of binomial filtering to the electric current at each time step to further reduce the noise. These strategies allowed us to evolve the system for an unprecedentedly long time of $\omega_{\mathrm{pi}}t=7000$, and up to $\omega_{\mathrm{pi}}t=12000$ in some cases. We initialize the upstream plasma with a velocity profile that is zero at the $x=0$ boundary, and linearly transitions to the upstream value $V_{0}$ at $x=100\,d_{\mathrm{i}}$ (Grošelj et al., 2024). The choice of initial flow profile suppresses the unphysically strong initial reflection of particles from the left wall before the shock is formed.
III Results
III.1 Shock Structure
@@
Interaction between the cosmic rays and the incoming upstream plasma results in a structure typical of the late stage of the Bell instability (Caprioli and Spitkovsky, 2014b; Crumley et al., 2019). In this paper, we define cosmic rays as nonthermal particles that are reflected and energized at the shock and propagate into the upstream. The quantitative definition and identification method of these cosmic rays in the simulations are discussed in Subsections III.2 and III.3. The fluctuations, particularly in the $B_{y}$ component (panel (c)), have a wavevector predominantly parallel to the ambient field, with a typical spatial scale of a few tens of ion skin depths, as expected for the Bell instability (Bell, 2004; Amato and Blasi, 2009). Locally, the fields reach near equipartition strength. The late nonlinear stage, which we are looking at, displays a large-scale structure of rarified cavities and dense filaments (panel (a)) and $B_{z}$ (panel (d)) (Caprioli and Spitkovsky, 2014a, b). Since the Bell modes, in their linear stage, have a circular polarization, the $B_{z}$ structure is similar to $B_{y}$, modulo this nonlinear density structure.
@@
These results are consistent with Crumley et al. (2019), in which the parameters in their fiducial run correspond to $\Gamma_{\mathrm{sh}}\sim 1.8,\theta_{B}=10^{\circ}$, and $\sigma=3\times 10^{-3}$, respectively, with our definitions, which is relatively similar to the run in Figure 2. Their simulation ends at $\omega_{\mathrm{pi}}t\sim 4000$. Since we have run the simulation longer than they have, the nonlinear filamentary structure is more pronounced.
@@
Panels (e,f) are the phase space densities $x-p_{x}$ of the two species, in which $p_{x}$ is normalized by the momentum of the upstream ions $m_{\mathrm{i}}\gamma_{0}v_{0}$. We can see that there is a significant amount of high-energy ions, with $|p_{x}|/m_{\mathrm{i}}\gamma_{0}v_{0}\gg 1$, in the downstream. However, there is a much smaller number of electrons in the same energy range. This implies that the particle acceleration efficiency is different between the two species. We will discuss particle acceleration in Subsection III.3.
@@
Figure 2: Snapshot of the $\sigma=10^{-3}$ shock, taken at $\omega_{\mathrm{pi}}t=7000$. Panel (a) is the plasma density $N_{\mathrm{e}}$ normalized by the upstream value. Panels (b-d) are the magnetic field components in units of the equipartition magnetic field, Equation (2). The colorbars are in a symmetric log scale, in which the range [0.01,0.01][-0.01,0.01] is in a linear scale, and values outside this range are in a log scale. Panels (e, f) are the phase space densities $f(x,p_{x})$ of ions and electrons, respectively. (An animation of this figure is available in the journal version.)
@@
Figure 3 shows the $\sigma=10^{-3.5}$ shock at the same time in the simulation $(\omega_{\mathrm{pi}}t=7000)$. The fluctuations that dominate in the upstream $B_{y}$ component, panel (c), still have a wave vector parallel to the ambient field, similarly to the $\sigma=10^{-3}$ case, but their amplitude is much smaller. This can be attributed to the cosmic ray current satisfying the high-current condition (Weidl et al., 2019). The high-current condition is satisfied when the cosmic ray current $J_{\mathrm{CR}}$ in the upstream becomes larger than a critical value, which depends on the upstream magnetization. We will discuss this in Subsection III.2. We can see in Figure 3 (e) that there are more reflected ions in the $\sigma=10^{-3.5}$ case, compared to the $\sigma=10^{-3}$ case (Figure 2 (e)). The difference in the number of downstream high-energy particles between ions and electrons is less significant than in the $\sigma=10^{-3}$ case.
@@
Figure 3: Snapshot of the $\sigma=10^{-3.5}$ shock, taken at $\omega_{\mathrm{pi}}t=7000$. The format is the same as in Figure 2. (An animation of this figure is available in the journal version.)
@@
Finally, Figure 4 is the $\sigma=10^{-4}$ shock at $\omega_{\mathrm{pi}}t=7000$. The magnetic field structure in this case is distinct from the previous cases. Most of the fluctuating magnetic field energy is in the $B_{z}$ component, panel (d), and little in $B_{y}$. The wavenumber vector is mostly perpendicular to the ambient field, i.e., $\bm{k}\perp\hat{\bm{B}}_{0}$, with a characteristic spatial scale of $d_{\mathrm{i}}k\sim 1$ in the downstream frame. These are the characteristics of the Weibel instability (Weibel, 1959; Fried, 1959). Note that cosmic rays are still streaming along the ambient field, and the magnetic field could alter the structure of the Weibel modes as compared to the case of an unmagnetized medium (Lemoine et al., 2014; Grassi et al., 2017). We discuss the unmagnetized $\sigma=0$ case in Appendix A.
@@
Figure 4: Snapshot of the $\sigma=10^{-4}$ shock, taken at $\omega_{\mathrm{pi}}t=7000$. The format is the same as in Figure 2. Compared to Figure 2, the shock structure is notably changed, as the precursor is now shaped by the Weibel instability. (An animation of this figure is available in the journal version.)
@@
Here, we have compared the structures of shocks with fixed $\Gamma_{\mathrm{sh}}\sim 2$ and $\theta_{B}=20^{\circ}$, varying the upstream $\sigma$. We have shown that the dominant physics in the upstream has a transition from the Bell instability at high magnetizations $\sigma\gtrsim 10^{-3}$ to the Weibel instability at lower magnetizations $\sigma\lesssim 10^{-4}$. In between, at $\sigma=10^{-3.5}$ we see a magnetic field structure, morphologically similar to Bell modes, but having a much smaller amplitude. In Subsection III.2, we discuss how the magnetization and cosmic ray properties determine the dominant instability. The $\sigma=10^{-4.5}$ case is not shown here because the overall structure is essentially the same as the $\sigma=10^{-4}$ case. Nevertheless, it will be included in the quantitative analysis of the following Subsections.
III.2 Cosmic Ray Driven Instabilities
@@
Here, we discuss the plasma streaming instabilities in the upstream, driven by the returning cosmic rays. In Subsection III.1, we discussed how the precursor is shaped by a competition between the Bell instability and the Weibel instability. It has recently been shown that the characteristics of Bell instability change when the current carried by the cosmic rays exceeds a certain threshold (Weidl et al., 2019). This is called the high-current regime. Lichko et al. (2025) suggests that the fluctuating fields $\delta B$ generated by Bell instability in the high-current regime saturate at a fraction of the ambient magnetic field energy. In their work, the typical saturation value was $\delta B/B_{0}\sim 5$. We find a similar value of $\delta B/B_{0}\sim 10$ for our parameters (see Appendix B). This is different from the classical low-current regime of Bell instability, in which a fixed fraction $\sim 0.1$ of the cosmic ray momentum flux is converted into the amplified fields (Caprioli and Spitkovsky, 2014b; Zacharegkas et al., 2024). The saturation energy density of the high-current Bell mode scales with $B_{0}^{2}$, while that of the low-current Bell mode scales with the cosmic ray momentum flux. It is essential to distinguish these two regimes, especially because we need to understand their competition with the Weibel instability. The high-current condition is satisfied when
@@
To investigate the nature of Bell mode saturation, we explicitly check Equation (3) in our PIC simulations. To this end, we define cosmic rays as the ions in the upstream region that have changed the sign of their $x$-momentum, at least once, i.e., that have been reflected back by the shock (Lemoine et al., 2019; Vanthieghem et al., 2022). We measure the cosmic ray ion four current $J^{\mu}_{\mathrm{CR}}$ at $x-x_{\mathrm{sh}}(t)]/d_{\mathrm{i}}=[200,300]$ in the simulation frame. Then we compute the upstream cosmic ray current along the background magnetic field as $J_{\mathrm{CR}}=J^{x}_{\mathrm{CR}}|_{\mathrm{u}}/\cos{\theta_{B}}$, where $|_{\mathrm{u}}$ indicates quantities measured in the upstream frame, and $\theta_{B}$ is defined in the upstream frame. In addition, we used the $y$ component and computed $J^{y}_{\mathrm{CR}}|_{\mathrm{u}}/\sin\theta_{B}$. The two methods give consistent results in this region. When the current is measured closer to the shock front, the two estimates are less consistent due to the change in the background plasma profile, e.g., resulting from the deceleration of the upstream flow (Vanthieghem et al., 2022). We note that we have assumed that the cosmic rays are dominated by ions. This is a common choice for Bell instability in non-relativistic shocks (Bell, 2004; Amato and Blasi, 2009). We confirm that this is also valid in trans-relativistic shocks, in Subsection III.3. The role of electron cosmic rays is discussed in Appendix C.
@@
Figure 5 shows the time evolution of the cosmic ray current averaged in the spatial interval $x-x_{\mathrm{sh}}(t)]/d_{\mathrm{i}}=[200,300]$. Panel (a) shows the raw $\eta$, whereas panel (b) corresponds to $\eta/\eta_{\mathrm{crit}}$. In panel (b), we can see that the $\sigma=10^{-3}$ shock eventually self-regulates the current to a subcritical value, consistent with the strong low-current Bell magnetic fluctuations in Figure 2. On the other hand, the cases with magnetization below $\sigma=10^{-3.5}$ remain supercritical, at least up to $\omega_{\mathrm{pi}}t=7000$. Looking at $1000\lesssim\omega_{\mathrm{pi}}t\lesssim 2000$ in panel (a), we notice that the two high-$\sigma$ cases are clearly distinct from the two low-$\sigma$ cases. This is presumably because of the difference in the ion reflection process. The classical reflection mechanism related to shock potential and mirror force (Balogh and Treumann, 2013; Burgess and Scholer, 2015) could apply to the high-$\sigma$ cases. On the other hand, the reflection in the low-$\sigma$ cases is governed by interaction with the small-scale Weibel fields in the vicinity of the shock front (Spitkovsky, 2008a; Kato and Takabe, 2008; Jikei and Amano, 2024; Parsons et al., 2024). This results in the $\sigma=10^{-3.5}$ case, Figure 3, being more supercritical, i.e., larger $\eta/\eta_{\mathrm{crit}}$, than the $\sigma=10^{-4}$ case, despite the larger $\eta_{\mathrm{crit}}$. Below $\sigma=10^{-4}$, the cosmic ray current is almost independent of $\sigma$; therefore, lower magnetization shocks would be more supercritical.